# Optimizing Large Language Models for ESG Activity Detection in Financial Texts

### v1

@@ Abstract
The integration of Environmental, Social, and Governance (ESG) factors into corporate decision-making is a fundamental aspect of sustainable finance. However, ensuring that business practices align with evolving regulatory frameworks remains a persistent challenge. AI-driven solutions for automatically assessing the alignment of sustainability reports and non-financial disclosures with specific ESG activities could greatly support this process. Yet, this task remains complex due to the limitations of general-purpose Large Language Models (LLMs) in domain-specific contexts and the scarcity of structured, high-quality datasets. In this paper, we investigate the ability of current-generation LLMs to identify text related to environmental activities. Furthermore, we demonstrate that their performance can be significantly enhanced through fine-tuning on a combination of original and synthetically generated data. To this end, we introduce ESG-Activities, a benchmark dataset containing 1,325 labeled text segments classified according to the EU ESG taxonomy. Our experimental results show that fine-tuning on ESG-Activities significantly enhances classification accuracy, with open models such as Llama 7B and Gemma 7B outperforming large proprietary solutions in specific configurations. These findings have important implications for financial analysts, policymakers, and AI researchers seeking to enhance ESG transparency and compliance through advanced natural language processing techniques.
Index Terms:
@@
Deep learning, Environmental management, Financial technology, Generative AI, Large Language Models, Machine learning, Natural language processing, Sustainability, Text classification.
I Introduction
@@
In recent years, driven by the widespread adoption of the Sustainable Development Goals (SDGs), the European Union has introduced principles and regulations aimed at helping organizations integrate environmental, social, and governance (ESG) factors into their operations and strategic decision-making. These initiatives encourage businesses and investors to assess and improve their environmental impact, fostering a more sustainable approach to economic activity [1]. As part of this effort, the ESG taxonomy [^1] https://finance.ec.europa.eu/sustainable-finance/tools-and-standards/eu-taxonomy-sustainable-activities_en has been established, offering a standardized framework to support sustainable practices across industries [2]. This resource enables companies to evaluate their activities in alignment with its criteria and report their performance in non-financial disclosures and sustainability reports.
However, aligning business practices with regulatory frameworks continues to be a challenge. Financial investors and other stakeholders often struggle to determine which aspects of corporate social responsibility correspond to specific activities regulated in the ESG taxonomy, making it difficult to accurately interpret and effectively utilize these reports. Indeed, a company can be associated with a vast amount of text that, in theory, can be mapped to various environmental activities described in the taxonomy. This text may include non-financial disclosures, marketing materials, website descriptions, product and service descriptions, corporate sustainability reports, and more. Thoroughly analyzing all this material is a complex and costly endeavour.
@@
The rise of social fintech technologies, which increasingly support intelligent solutions for social impact investing, presents a promising opportunity to leverage advanced computational techniques for analyzing financial texts and extracting sustainability-related insights. For instance, several AI models have been developed to classify financial documents based on relevant SDGs [3, 4]. The rapid advancement of large language models (LLMs) in recent years has further improved automated solutions for the analysis of financial and sustainability-related texts [5, 6, 7, 8, 9]. However, categorizing text based on specific environmental activities, such as those defined in the ESG taxonomy, remains a significant challenge. Examples of these activities include promoting low-carbon rail transport, ensuring green port operations with zero direct CO2 emissions, or expanding infrastructure for cycling and personal electric mobility. This task is considerably more complex than linking text to broad SDG categories, as it requires a fine-grained understanding of the actions and activities described in a document. General-purpose LLMs struggle with this level of specificity. Overcoming these limitations, therefore, requires fine-tuning such models on high-quality, domain-specific datasets – an effort often hindered by the scarcity of training data in this domain.
@@
To fine-tune and evaluate a range of models on their task, we introduce ESG-Activities, a novel benchmark consisting of 1,325 texts classified according to ESG activities. Since we intended assess the capability of synthetic data in supporting AI systems in this domain, the training set of ESG-Activities includes a mixture of manually curated data from human experts and synthetic data generated by LLMs. In contrast, the test set consists exclusively of items curated by human experts to ensure a high-quality evaluation.
The evaluation demonstrates that fine-tuning the system with a combination of manually crafted and synthetically generated data leads to significant improvements, not only compared to zero-shot learning but also over models fine-tuned exclusively on the manually annotated dataset. Furthermore, we show that relatively small open-source models, such as Llama 7B, can achieve excellent performance when fine-tuned on ESG-Activities, even surpassing large proprietary models in certain configurations.
@@
  • We provide the complete codebase for our analysis to ensure reproducibility [^3] Codebase - https://github.com/Mattia-Brt/Fine_tuning_LLM.
@@
The remainder of this paper is structured as follows. Section II discusses the use case and provides a formal definition of the task. Section III introduces the new benchmark. Section IV describes the methodology adopted in our analysis. Section V presents the evaluation and discusses key insights. Section VI reviews related work. Finally, Section VII concludes the paper and outlines future research directions.
II Task definition and use case
II-A The role of ESG taxonomy for social fintech companies
@@
In the European landscape, the analysis of social and environmental impact investments and the assessment of the sustainability of companies, including crowdfunding platforms or fintech, are becoming increasingly crucial. In 2021, the European Commission defined a set of ambitious policies related to the so-called “European Green Deal” with the overarching goal of making Europe climate-neutral by 2050. Among other initiatives, the European Green Deal has established a taxonomy that defines the activities companies must undertake to enhance their contributions to ESG aspects. This taxonomy is becoming a crucial benchmark for European investors, who utilize it to assess the sustainability of their investment decisions. Investors require a thorough understanding of the extent to which companies comply with the European ESG taxonomy. This is essential for making informed investment decisions that align with their values and contribute to sustainable development. However, the complexity of the ESG taxonomy and the vast amount of financial documentation can make it challenging for investors to independently assess a company’s compliance. As a result, there is an increasing need for tools that can streamline this process by automatically identifying and annotating relevant sections within financial documents. These tools represent sustainable applications of Social Fintech, showcasing how AI can enhance corporate social responsibility. By enabling investors to identify companies genuinely committed to sustainability, these technologies help direct capital toward initiatives aligned with SDGs, such as those outlined in the European Green Deal.
@@
Furthermore, semi-automatic annotation tools can be seamlessly integrated into Social Fintech platforms, including crowdfunding platforms, reinforcing their dedication to sustainability and social impact. Automating ESG analysis with AI enhances transparency and promotes responsible investment opportunities, ultimately fostering a more sustainable and equitable financial system.
II-B Task definition
@@
In this paper, we describe the process of developing an LLM tailored to verify if a text portion from NFD relates to a particular item in the ESG taxonomy. The task is framed as a binary classification problem, where the goal is to train the LLM to answer the question: Does c pertain to i?. Specifically, for a text segment $c$ and an item $i$ – in this case, the description of an ESG activity – the formulated fine-tuned LLM function $f(c,i)$ is expressed as:
@@
The aim of the fine-tuned LLM is to learn a mapping function $f$ that accurately assigns each pair $(c,i)$ to the correct class.
@@
The fine-tuning process was performed using the Low-Rank Adaptation (LoRA) technique [10], which updates only a low-rank subspace of the model’s pre-trained weight matrices, reducing the number of trainable parameters while maintaining efficiency. Let $W$ be a weight matrix from the pre-trained model. During fine-tuning, the adaptation introduces a low-rank update $\Delta W$:
$\Delta W=AB,$
where $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, $r\ll\min(m,n)$, ensuring efficiency in storage and computation.
The updated weight matrix becomes:
$W^{\prime}=W+\Delta W=W+AB.$
By freezing the original weights $W$ and only optimizing $A$ and $B$, LoRA enables efficient fine-tuning with minimal computational overhead.
III The ESG-Activities Benchmark
@@
Since the original descriptions were too long, we generated shorter versions using GPT-4 as a rephraser. We then indexed the NFDs in a vector database and employed a Retrieval-Augmented Generation (RAG) [11] pipeline to retrieve the most relevant text chunks for each activity description. This process yielded 265 candidate mappings between text segments and ESG actions.
@@
Since we aimed to verify the effectiveness of syntactic data in this domain, we then expanded the training set with artificial data. Specifically, we used ChatGPT-4o to generate five alternative formulations for each of the original 212 sentences in the training set, ensuring they conveyed the same meaning while using different wording. This process resulted in 1,060 additional sentences, which were labelled as ’synthetic data’ and incorporated into the training set. This setup allows users of the benchmark to fine-tune their models using only the original data (212 items) or a combination of original and syntactic data (1272 items).
IV Methodology
@@
IV-A System Architecture
The study presented in this paper is part of a broader project aimed at developing ESGQuest, a research prototype designed to support the annotation of NFD with ESG-related activities. This system is built on a RAG pipeline. Given a document about a company and a set of relevant NACE codes, ESGQuest aims to identify text segments that align with ESG activities associated with the specified NACE codes. The results are presented as an annotated PDF, which can be further refined by human users. [^4] A demo of this system is available at https://esgquest.datai.disco.unimib.it. The ESGQuest pipeline operates as follows.
First, ESGQuest selects ESG activities based on the given NACE codes using a predefined mapping schema. NACE codes (Nomenclature statistique des Activités économiques dans la Communauté Européenne) are a European industry standard classification system that groups organizations based on their economic activities.555 These codes are used in the EU Taxonomy to assess the environmental sustainability of economic activities. For more details, visit https://ec.europa.eu/eurostat/web/nace This step is optional but generally beneficial, as many ESG activities described in the taxonomy are relevant only to specific industrial sectors (e.g., transport, energy). Including irrelevant activities in the analysis would not only waste computational resources but also introduce unnecessary complexity and potential confusion.
@@ IV-A System Architecture
IV-B Large Language Models
@@
Llama, introduced by Hugo Touvron et al. in 2023 [5], is a family of decoder-only language models designed to deliver high performance and computational efficiency. The Llama models range from 1 to 90 billion parameters, making them suitable for a wide variety of natural language processing applications. In this study, we evaluate three versions: Llama 3B, Llama 2 7B, and Llama 3 8B, all available on official platforms such as Hugging Face.
Llama 3B [^7] https://huggingface.co/openlm-research/open_Llama_3b is an open-source model with 3 billion parameters. The model incorporates architectural enhancements such as pre-normalization and the SwiGLU activation function [12], which contribute to training stability and efficiency.
Llama 2 7B [^8] https://huggingface.co/meta-Llama/Llama-2-7b-hf, represents an evolution of the first generation of Llama [13]. It was trained using additional data and optimized techniques to improve generalization. This model excels in zero-shot and few-shot learning tasks and directly competes with larger models like GPT-3 and Chinchilla [14].
Llama 3 8B [^9] https://huggingface.co/meta-llama/Meta-Llama-3-8B, is the latest iteration in the series. This model, featuring 8 billion parameters, was trained on a dataset exceeding 15 trillion tokens [15]. It incorporates significant advancements such as odds ratio preference optimization [16], which optimizes performance and alignment.
The Llama series was trained on diverse datasets, comprising publicly available data such as CommonCrawl, C4 [17], GitHub, Wikipedia, Books3 [18], ArXiv, and Stack Exchange.
These models have demonstrated excellent performance in benchmarks such as NaturalQuestions [19] and TriviaQA [20], often surpassing larger models like Gopher (280B) [21] and Chinchilla (70B) [14]. For instance, the Llama 33B model outperformed both Gopher and Chinchilla in question-answering tasks on the NaturalQuestions benchmark, demonstrating that the Llama architecture effectively balances size and performance.
IV-B2 Gemma.
@@
Introduced by Google in 2024, Gemma [6] is a family of open-source language models derived from the Gemini series [7], designed to achieve an optimal balance between computational efficiency and performance across various natural language processing tasks. The models are available in three configurations – 2B, 7B, and 27B parameters – each tailored for specific applications. Gemma employs a decoder-only architecture with advanced features such as multi-query attention to reduce memory overhead [22], Rotary Position Embeddings (RoPE) [17] to improve positional representation, GeGLU activations for better convergence during training, and RMSNorm to enhance stability on large-scale datasets. The models were trained on diverse publicly available data, including CommonCrawl, C4 [17], GitHub repositories, ArXiv papers, Wikipedia, and Books3 [18], ensuring broad linguistic coverage and domain-specific expertise. The 2B model [^10] https://huggingface.co/google/gemma-2b is optimized for on-device deployments and CPU-based inference, making it efficient for general-purpose applications. The 7B variant [^11] https://huggingface.co/google/gemma-7b is fine-tuned for complex tasks like advanced coding, scientific reasoning, and mathematical problem-solving, outperforming comparable models such as Llama 7B [5] and often rivalling Mistral 7B [8]. The Recurrent Gemma 2B model [^12] https://huggingface.co/google/recurrentgemma-2b introduces a recurrent mechanism that enhances performance on sequential reasoning tasks and long-context processing, making it particularly effective for multi-turn dialogues and structured data analysis [23]. Across multiple benchmarks, including coding, reasoning, and question answering, Gemma models have demonstrated competitive performance.
IV-B3 Mistral.
@@
Mistral, introduced by Albert Q. Jiang et al. in 2023 [8], is a highly efficient decoder-only language model designed for a wide array of natural language processing tasks, excelling in areas such as reasoning, mathematics, code generation, and commonsense reasoning. With 7.3 billion parameters, Mistral was specifically engineered to achieve a balance between computational efficiency and performance. The model incorporates several advanced techniques that make it highly effective for tasks that demand long-context processing and memory optimization. One such technique is the sliding window attention (SWA) [24], which enables the model to process long sequences of input in overlapping chunks, allowing it to handle inputs of arbitrary length with minimal computational overhead. In addition, Mistral utilizes grouped-query attention (GQA) [25], a method that optimizes inference speed by reducing memory usage during the attention computation phase.
Mistral’s ability to maintain performance despite having a smaller parameter count than other models makes it a powerful tool for practical applications, especially where computational resources may be limited. It consistently outperforms larger models like Llama 2 13B [5], Llama 1 34B [26], and other established benchmarks across a variety of metrics. This is especially evident in reasoning-intensive tasks, where Mistral exhibits superior generalization capabilities.
The 7B variant of Mistral [^13] https://huggingface.co/mistralai/Mistral-7B-v0.1 has been fine-tuned to support diverse use cases, including scientific research, where accurate interpretation and generation of complex texts are critical. The model is also capable of achieving state-of-the-art results on tasks requiring a deep understanding of specialized knowledge, such as those found in academic papers or technical documents.
IV-B4 Gpt-4o Mini.
@@
GPT-4o Mini is a decoder-only language model developed by OpenAI as a faster and more cost-effective alternative to GPT-4o. GPT-4o is a comprehensive multimodal model capable of processing and generating text, images, and audio, designed for complex tasks requiring high computational power and precision. It features a 128K-token context window and supports up to 16K output tokens per request, with a knowledge cutoff in October 2023 [9]. Despite its smaller size, GPT-4o Mini outperforms previous models, such as GPT-3.5 Turbo, in academic benchmarks, demonstrating superior performance in textual intelligence and multimodal reasoning [27]. It retains the same context window and output token limits as GPT-4o, ensuring robust performance across various applications. Additionally, GPT-4o Mini incorporates an instruction hierarchy method to enhance resistance to jailbreaks and prompt injections, improving security for deployment in diverse settings [28].
IV-C Experimental Setup

### v2

@@ Abstract.
The integration of Environmental, Social, and Governance (ESG) factors into corporate decision-making is a fundamental aspect of sustainable finance. However, ensuring that business practices align with evolving regulatory frameworks remains a persistent challenge. AI-driven solutions for automatically assessing the alignment of sustainability reports and non-financial disclosures with specific ESG activities could greatly support this process. Yet, this task remains complex due to the limitations of general-purpose Large Language Models (LLMs) in domain-specific contexts and the scarcity of structured, high-quality datasets. In this paper, we investigate the ability of current-generation LLMs to identify text related to environmental activities. Furthermore, we demonstrate that their performance can be significantly enhanced through fine-tuning on a combination of original and synthetically generated data. To this end, we introduce ESG-Activities, a benchmark dataset containing 1,325 labeled text segments classified according to the EU ESG taxonomy. Our experimental results show that fine-tuning on ESG-Activities significantly enhances classification accuracy, with open models such as Llama_7B and Gemma_7B outperforming large proprietary solutions in specific configurations. These findings have important implications for financial analysts, policymakers, and AI researchers seeking to enhance ESG transparency and compliance through natural language processing techniques.
@@
Deep learning, Environmental management, Financial technology, Generative AI, Large Language Models, Machine learning, Natural language processing, Sustainability, Text classification
copyright: acmlicensedjournalyear: 2025conference: 6th ACM International Conference on AI in Finance; November 15–18, 2025; Singapore, Singaporebooktitle: 6th ACM International Conference on AI in Finance (ICAIF ’25), November 15–18, 2025, Singapore, Singaporedoi: 10.1145/3768292.3770371isbn: 979-8-4007-2220-2/2025/11ccs: Theory of computation Models of learningccs: Theory of computation Data modelingccs: Theory of computation Data integrationccs: Computing methodologies Natural language generationccs: Computing methodologies Natural language processingccs: Computing methodologies Cross-validation
1. Introduction
@@
In recent years, driven by the widespread adoption of the Sustainable Development Goals (SDGs), the European Union has introduced principles and regulations aimed at helping organizations integrate environmental, social, and governance (ESG) factors into their operations and strategic decision-making. These initiatives encourage businesses and investors to assess and improve their environmental impact, fostering a more sustainable approach to economic activity (kostetckaia2022sustainable). As part of this effort, the ESG taxonomy [^1] https://finance.ec.europa.eu/sustainable-finance/tools-and-standards/eu-taxonomy-sustainable-activities_en has been established, offering a standardized framework to support sustainable practices across industries (dumrose2022disaggregating). This resource enables companies to evaluate their activities in alignment with its criteria and report their performance in non-financial disclosures and sustainability reports.
@@
However, aligning business practices with regulatory frameworks continues to be a challenge. Financial investors and other stakeholders often struggle to determine which aspects of corporate social responsibility correspond to specific activities regulated in the ESG taxonomy, making it difficult to accurately interpret and effectively utilize these reports. Indeed, a company can be associated with a vast amount of text that, in theory, can be mapped to various environmental activities described in the taxonomy. This text may include non-financial disclosures, marketing materials, website descriptions, product and service descriptions, corporate sustainability reports, and more. The rise of social fintech technologies, which increasingly support intelligent solutions for social impact investing, presents a promising opportunity to leverage advanced computational techniques for analyzing financial texts and extracting sustainability-related insights. For instance, several AI models have been developed to classify financial documents based on relevant SDGs (guariso2023automatic; hajikhani2022mapping). The rapid advancement of large language models (LLMs) in recent years has further improved automated solutions for the analysis of financial and sustainability-related texts (llama; gemma; gemini; mistral; openai2023gpt4o). However, categorizing text based on specific environmental activities, such as those defined in the ESG taxonomy, remains a significant challenge such as promoting low-carbon rail transport, ensuring green port operations with zero direct CO2 emissions. This task is considerably more complex than linking text to broad SDG categories, as it requires a fine-grained understanding of the actions and activities described in a document. General-purpose LLMs struggle with this level of specificity. Overcoming these limitations, therefore, requires fine-tuning such models on high-quality, domain-specific datasets – an effort often hindered by the scarcity of training data in this domain.
@@
To fine-tune and evaluate a range of models on their task, we introduce ESG-Activities, a novel benchmark consisting of 1,325 texts classified according to ESG activities. Since we intended to assess the capability of synthetic data in supporting AI systems in this domain, the training set of ESG-Activities includes a mixture of manually curated data from human experts and synthetic data generated by LLMs. In contrast, the test set consists exclusively of items curated by human experts to ensure a high-quality evaluation.
In the evaluation, we simulate realistic low-resource scenarios and therefore focus primarily on relatively lightweight LLMs with sizes between 2B and 7B, drawn from several families including LLaMA, Gemma, and Mistral. We also compare these solutions with alternative approaches, such as the proprietary GPT4o-mini and the domain-specific, encoder-only ESG-BERT.
The experiments demonstrate that fine-tuning the system with a combination of manually crafted and synthetically generated data leads to significant improvements, not only compared to zero-shot learning but also over models fine-tuned exclusively on the manually annotated dataset. Furthermore, we show that relatively small open-source models, such as Llama 7B, can achieve excellent performance when fine-tuned on ESG-Activities, even surpassing large proprietary models in certain configurations.
@@
  • We provide the complete codebase for our analysis to ensure reproducibility [^3] Codebase - https://github.com/Mattia-Brt/Fine_tuning_LLM/tree/main.
@@
The remainder of this paper is structured as follows. Section 2 discusses the use case and provides a formal definition of the task. Section 3 introduces the new benchmark. Section 4 describes the methodology adopted in our analysis. Section 5 presents the evaluation and discusses key insights. Section 6 reviews related work. Finally, Section 7 concludes the paper and outlines future research directions.
2. Task definition and use case
2.1. The role of ESG taxonomy for Social Fintech
@@
In the European landscape, the analysis of social and environmental impact investments and the assessment of the sustainability of companies, including crowdfunding platforms or fintech, are becoming increasingly crucial. In 2021, the European Commission defined a set of ambitious policies related to the so-called “European Green Deal” with the overarching goal of making Europe climate-neutral by 2050. Among other initiatives, the European Green Deal has established a taxonomy that defines the activities companies must undertake to enhance their contributions to ESG aspects. This taxonomy is becoming a crucial benchmark for European investors, who utilize it to assess the sustainability of their investment decisions. Investors require a thorough understanding of the extent to which companies comply with the European ESG taxonomy. This is essential for making informed investment decisions that align with their values and contribute to sustainable development. However, the complexity of the ESG taxonomy and the vast amount of financial documentation can make it challenging for investors to independently assess a company’s compliance. As a result, there is an increasing need for tools that can streamline this process by automatically identifying and annotating relevant sections within financial documents. By enabling investors to identify companies genuinely committed to sustainability, these technologies help direct capital toward initiatives aligned with SDGs, such as those outlined in the European Green Deal.
@@
Automating ESG analysis with AI enhances transparency and promotes responsible investment opportunities, ultimately fostering a more sustainable and equitable financial system.
2.2. Task definition
@@
In this paper, we describe the process of developing a language model tailored to verify if a text portion from NFD relates to a particular item in the ESG taxonomy. The task is framed as a binary classification problem, where the goal is to train the model to answer the question: Does c pertain to i? Specifically, for a text segment $c$ and an item $i$ – in this case, the description of an ESG activity – the model implements the function $f(c,i)$ defined as:
@@
The fine-tuning process was performed using the Low-Rank Adaptation (LoRA) technique (heidloff2023efficient), which updates only a low-rank subspace of the model’s pre-trained weight matrices, reducing the number of trainable parameters while maintaining efficiency.
3. The ESG-Activities Benchmark
@@
Since the original descriptions were very long, we generated shorter versions using GPT-4 as a rephraser. We then indexed the NFDs in a vector database and employed a Retrieval-Augmented Generation (RAG) (lewis2020retrieval) pipeline to retrieve the most relevant text chunks for each activity description. This process yielded 265 candidate mappings between text segments and ESG actions.
@@
Since we aimed to verify the effectiveness of synthetic data in this domain, we then expanded the training set with artificial data. Specifically, we used ChatGPT-4o to generate five alternative formulations for each of the original 212 sentences in the training set, ensuring they conveyed the same meaning while using different wording. This process resulted in 1,060 additional sentences, which were labelled as ’synthetic data’ and incorporated into the training set. This setup allows users of the benchmark to fine-tune their models using only the original data (212 items) or a combination of original and synthetic data (1272 items).
4. Methodology
@@
4.1. System Architecture
The study presented in this paper is part of a broader project aimed at developing ESGQuest, a research prototype designed to support the annotation of NFDs with ESG-related activities. This system is built on a RAG pipeline. Given a document about a company and a set of relevant NACE codes, ESGQuest aims to identify text segments that align with ESG activities associated with the specified NACE codes. The results are presented as an annotated PDF, which can be further refined by human users. [^4] A demo of this system is available at https://esgquest.datai.disco.unimib.it. The ESGQuest pipeline operates as follows.
First, ESGQuest selects ESG activities based on the given NACE codes using a predefined mapping schema. NACE codes (Nomenclature statistique des Activités économiques dans la Communauté Européenne) are a European industry standard classification system that groups organizations based on their economic activities. [^5] These codes are used in the EU Taxonomy to assess the environmental sustainability of economic activities. For more details, visit https://ec.europa.eu/eurostat/web/nace This step is optional but generally beneficial, as many ESG activities described in the taxonomy are relevant only to specific industrial sectors (e.g., transport, energy). Including irrelevant activities in the analysis would not only waste computational resources but also introduce unnecessary complexity and potential confusion.
@@ 4.1. System Architecture
4.2. Large Language Models
@@
Llama, introduced by Hugo Touvron et al. in 2023 (llama), is a family of decoder-only language models designed to achieve both high performance and computational efficiency. The Llama models span from 1 to 90 billion parameters, making them suitable for a wide range of natural language processing tasks. They were trained on diverse datasets, including publicly available sources such as CommonCrawl, C4 (t5), GitHub, Wikipedia, Books3 (books3), ArXiv, and Stack Exchange. In this study, we evaluate three models from the Llama family: Llama 3B [^7] https://huggingface.co/openlm-research/open_Llama_3b, Llama 2 7B [^8] https://huggingface.co/meta-Llama/Llama-2-7b-hf, and Llama 3 8B [^9] https://huggingface.co/meta-llama/Meta-Llama-3-8B.
4.2.2. Gemma.
@@
Introduced by Google in 2024, Gemma (gemma) is a family of open-source language models derived from the Gemini series (gemini), designed to achieve an optimal balance between computational efficiency and performance across various natural language processing tasks. Gemma employs a decoder-only architecture with advanced features such as multi-query attention to reduce memory overhead (team2024gemma), Rotary Position Embeddings (RoPE) (t5) to improve positional representation, GeGLU activations for better convergence during training, and RMSNorm to enhance stability on large-scale datasets. The models were trained on diverse publicly available data, including CommonCrawl, C4 (t5), GitHub repositories, ArXiv papers, Wikipedia, and Books3 (books3), ensuring broad linguistic coverage and domain-specific expertise. In this study, we evaluate three models from the Gemma family: Gemma_2B [^10] https://huggingface.co/google/gemma-2b, Gemma_7B [^11] https://huggingface.co/google/gemma-7b, and RecurrentGemma_2B [^12] https://huggingface.co/google/recurrentgemma-2b.
4.2.3. Mistral.
@@
Mistral, introduced by Albert Q. Jiang et al. in 2023 (mistral), is a highly efficient decoder-only language model designed for a wide array of natural language processing tasks, excelling in areas such as reasoning, mathematics, code generation, and commonsense reasoning. With 7.3 billion parameters, Mistral was specifically engineered to achieve a balance between computational efficiency and performance. The model incorporates several advanced techniques that make it highly effective for tasks that demand long-context processing and memory optimization. One such technique is the sliding window attention (SWA) (swa), which enables the model to process long sequences of input in overlapping chunks, allowing it to handle inputs of arbitrary length with minimal computational overhead. In this study, we employ the 7B variant of Mistral [^13] https://huggingface.co/mistralai/Mistral-7B-v0.1, which has been fine-tuned to support a variety of use cases, including scientific research where the accurate interpretation and generation of complex texts are essential.
4.2.4. Gpt-4o Mini.
@@
GPT-4o Mini is a decoder-only language model developed by OpenAI as a faster and more cost-effective alternative to GPT-4o. GPT-4o is a comprehensive multimodal model capable of processing and generating text, images, and audio, designed for complex tasks requiring high computational power and precision. It features a 128K-token context window and supports up to 16K output tokens per request, with a knowledge cutoff in October 2023 (openai2023gpt4o). Despite its smaller size, GPT-4o Mini outperforms previous models, such as GPT-3.5 Turbo, in academic benchmarks, demonstrating superior performance in textual intelligence and multimodal reasoning (openai2023gpt4omini). It retains the same context window and output token limits as GPT-4o, ensuring robust performance across various applications. Additionally, GPT-4o Mini incorporates an instruction hierarchy method to enhance resistance to jailbreaks and prompt injections, improving security for deployment in diverse settings (openai2023gpt4osafety).
4.2.5. ESG-BERT
ESG-BERT (nbroad_ESG-BERT) is a domain-specific variant of BERT, pre-trained on a large corpus of sustainability-related texts. It is designed to capture the terminology and linguistic nuances of the ESG domain, and it achieves strong performance on text classification and sentiment analysis tasks in this context. Specifically, ESG-BERT has been shown to effectively identify key concepts related to sustainability and corporate social responsibility. In this study, we include ESG-BERT as a baseline for comparison with fine-tuned LLMs.
4.3. Experimental Setup